OpenStudy (anonymous):

will fan and medal, i really need help on this! A student reasons that (4, 1) is a solution of y = 3x + 1 because x = 1 when y = 4. Explain and correct the student’s error.

OpenStudy (anonymous):

i really need an explanation on how to do it

OpenStudy (texaschic101):

(4,1) = (x,y) x is not 1...x is 4 y is not 4...y is 1

OpenStudy (anonymous):

Thank you, im confused as to how i didn't see that haha :)

OpenStudy (anonymous):

wait is that correct tho ..@texaschic101

OpenStudy (texaschic101):

oh..wait....the solution set is written wrong..it should be (1,4) y = 3x + 1 4 = 3(1) + 1 4 = 3 + 1 4 = 4 (correct)

OpenStudy (anonymous):

nvm nvm nvm i see how

OpenStudy (texaschic101):

the correct solution set is (1,4) not (4,1)